### v2

# Does LLM Focus on the Right Words? Mitigating Context Bias in LLM-based Recommenders

Abstract.
Large language models (LLMs), owing to their extensive open-domain knowledge and semantic reasoning capabilities, have been increasingly integrated into recommender systems (RS). However, a substantial gap remains between the pre-training objectives of LLMs and the specific requirements of recommendation tasks. To address this gap, supervised fine-tuning (SFT) is commonly performed on specially curated recommendation datasets to further enhance their predictive ability. Despite its success, SFT exhibits a critical limitation: it induces Context Bias, whereby the model over-relies on auxiliary tokens—such as task descriptions and prefix-generated tokens—while underutilizing core user interaction tokens that encode user-specific preferences. This bias not only undermines recommendation accuracy but also raises unfairness concerns.
To address this issue, we propose Group Distributionally Robust Optimization-based Tuning (GDRT), a novel fine-tuning paradigm that enforces consistent model performance across token groups with varying degrees of relevance to auxiliary tokens. By adaptively upweighting underperforming groups, typically those weakly correlated with auxiliary tokens, GDRT shifts the model’s attention from superficial auxiliary cues to informative user interaction tokens, thereby mitigating context bias. Extensive experiments conducted on three public datasets demonstrate that GDRT effectively mitigates context bias, yielding substantial improvements in recommendation accuracy (with an average NDCG@10 gain of 24.29%) and significantly enhancing recommendation fairness. The code is available at https://github.com/WANGBohaO-jpg/GDRT.
Sequential Recommendation; Large Language Model; Bias
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
1. Introduction
With remarkable open-domain knowledge and semantic reasoning capabilities (Guo et al., 2025; Achiam et al., 2023), Large Language Models (LLMs) have been extensively explored for integration into recommendation systems (RS) (Wu et al., 2024). One prominent approach involves positioning LLMs as the central recommendation backbone (Li et al., 2023c; Bao et al., 2025, 2023; Zhu et al., 2024; Wang et al., 2025b, 2024b; Liao et al., 2024b; Kim et al., 2024; Hou et al., 2024; Wang and Lim, 2023; Tan et al., 2024; Zheng et al., 2024a). These methods express items as textual descriptions (e.g., titles), and construct language prompts based on users’ past interactions, which are then used to instruct LLMs to predict users’ future interactions. Figure 1 illustrates the mechanism of such LLM-based recommendation. LLMs operate at a fine-grained semantic token level, sequentially generating tokens of the predicted items by analyzing their nuanced semantic relations with the user’s previous interactions and other auxiliary information (e.g., task descriptions, prefix tokens of the prediction). This fine-grained paradigm enables the capture of subtle semantic patterns in user interests and thus represents a promising direction for advancing recommender systems (Bao et al., 2025).
Figure 1. Illustration of LLM‑based recommendations and context bias, wherein the model exhibits an over‑reliance on auxiliary tokens and insufficient utilization of interaction tokens during generation.
Figure 2. Ratio of attribution values between auxiliary tokens and user interaction tokens before and after SFT. Left: task description vs. user interaction tokens. Right: prefix tokens of predicted item (take the first token) vs. user interaction tokens.
To better align LLMs with recommendation objectives and capture collaborative filtering signals, Supervised Fine-Tuning (SFT) is commonly employed (Bao et al., 2025, 2023; Liao et al., 2024b; Zhu et al., 2024). In this strategy, each prompt is paired with the target item description, and the LLM is fine-tuned to generate the correct prediction. This procedure enables the model to capture semantic relationships between prompts and targets present in the training data, often resulting in substantial performance gains.
However, we find that SFT introduces a significant Context Bias. Specifically, SFT drives the model to over-rely on auxiliary tokens (e.g., task descriptions or prefix tokens) while under-utilizing core interaction tokens that encode user personalized preferences as shown in Figure 1. To verify this, we conduct Feature Ablation Attribution analysis (Kokhlikyan et al., 2020; Miglani et al., 2023), a standard approach to quantify each token’s contribution to model predictions. Figure 2 shows that SFT dramatically amplifies the relative impact of auxiliary tokens while significantly suppressing the influence of interaction tokens, with the influence ratio shifting from about 1:1 before fine-tuning to more than 6:1 afterward as measured on typical Amazon datasets.
Figure 3. Distribution of Top‑1 recommended items generated by SFT-trained LLMs across five group defined according to the semantic relevance of items to the auxiliary tokens (Group 1: highest relevance, Group 5: lowest relevance). We also present the distribution of the test set across the same five groups for comparison.
This over-reliance reveals shortcut learning: LLMs simply memorize correlations with frequently occurring auxiliary tokens rather than grounding predictions in user-specific preferences. Such bias not only undermines recommendation accuracy but also raises serious fairness concerns. Specifically, this bias skews recommendations toward a narrow subset of items whose tokens exhibit higher semantic relevance to auxiliary tokens. To verify this, we divided items into five groups based on their relevance to auxiliary tokens. As shown in Figure 3, over 80% of recommended items fall into the highest relevance group, while this group comprises only 20% of target items in the test set. These findings motivate our core research question: How can we mitigate context bias in LLM-based recommenders?
To address this, we propose a novel fine-tuning strategy GDRT, that leverages Group Distributionally Robust Optimization (Group DRO) (Sagawa et al., 2019) to mitigate context bias. We first group training instances according to the semantic relevance between the target token and the auxiliary tokens, which can be evaluated by the LLM’s predictive probability when user history is masked. Group DRO then enforces LLMs to perform consistently well across all groups, regardless of their relevance strength with auxiliary tokens. This optimization objective naturally shifts the LLM’s attention away from auxiliary tokens toward user-specific interaction tokens, as simple reliance on shortcut auxiliary tokens results in poor performance on groups with weaker correlations. Importantly, GDRT is easy to implement and computationally efficient, requiring only high‑efficiency group construction and dynamic sample weighting during training. It can be seamlessly integrated into various LLM-based recommenders, yielding improvements in both accuracy and fairness.
Our main contributions are:
  • We provide a comprehensive empirical analysis revealing that SFT in LLM-based recommendation induces significant context bias, negatively affecting both accuracy and fairness.
  • We propose GDRT, a Group DRO-based fine-tuning strategy, to effectively mitigate context bias in LLM-based recommendation.
  • We conduct extensive experiments demonstrating that GDRT achieves state-of-the-art recommendation performance in both accuracy and fairness metrics.
2. Preliminary
2.1. LLM-based Recommendation
Following previous work (Liao et al., 2024b; Bao et al., 2025; Na et al., 2024; Bao et al., 2023; Lin et al., 2024; Zheng et al., 2024b), this paper also focuses on sequential recommendation (Kang and McAuley, 2018), a conventional recommendation scenario in practice. Let $\mathcal{V}$ denote the set of items in the recommendation system. Given a user’s historical interaction sequence $S=\{s_{1},s_{2},...,s_{n}\}$, where each $s_{i}\in\mathcal{V}$ represents the $i$-th interacted item, the goal of the RS is to predict the user’s next interaction $s_{n+1}$ that the user is likely to interact with.
The remarkable success of LLMs across diverse domains (Wu et al., 2023b; Ouyang et al., 2022; Tang et al., 2024; Chen et al., 2025c; Wang et al., 2025a; Chen et al., 2025b) has spurred growing interest in their application to recommendation systems (RS) (Wu et al., 2024). A prominent approach is to directly leverage LLMs as recommenders (Li et al., 2023c). As shown in Figure 1, this paradigm constructs a language prompt $x=[\,x^{\text{task}};\,x^{\text{user}}\,]$, where $x^{\text{task}}$ represents the task description and $x^{\text{user}}$ denotes the textual form (e.g., titles) of a user’s historical interactions. This prompt then guides the LLM to generate the descriptions of recommended items $y$. Notably, LLMs operate at a fine-grained semantic token level, sequentially generating tokens of the predicted items according to the model estimated probability $P_{\theta}(y_{t}|x,y_{<t})$, where $y_{t}$ denotes the $t$-th predictive tokens and $y_{<t}$ denotes the prefix tokens of the prediction. This fine-grained token-level paradigm has the potential to capture subtle semantic patterns in user preferences.
To align LLMs with recommendation objectives, supervised fine-tuning (SFT) is commonly applied, fine-tuning all or part of the model parameters using recommendation data (Bao et al., 2025, 2023; Liao et al., 2024b; Zhu et al., 2024). In this process, the training data is reorganized into a set of prompt–target pairs $\mathcal{D}=\{(x_{i},y^{*}_{i})\}_{i=1}^{N}$ where each $x_{i}$ is the constructed prompt and $y^{*}_{i}$ is the textual description of the target item. The LLM is optimized with the following log-likelihood objective:
(1) $\mathcal{L}_{SFT}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{|y^{*}_{i}|}\log P_{\theta}\!\left(y^{*}_{i,t}\mid x_{i},y^{*}_{i,<t}\right)$
where $|y^{*}_{i}|$ denotes the token length of the target item description. SFT increases the generative probability of the target item, encouraging the LLM to capture the inherent token-level semantic correlations between each target token $y_{i,t}^{*}$ and the user interactions $x^{\text{user}}_{i}$, task descriptions $x^{\text{task}}$ and the prefix tokens $y^{*}_{i,<t}$. This process often yields substantial performance improvements (Bao et al., 2023).
2.2. Analyses on Context Bias
In this section, we first identify the context bias in fine-tuning LLM for recommendation, followed by discussing its negative effect. We then analyze the underlying causes of this bias and discuss why existing methods can not effectively address this issue.  [^1] The configuration of the empirical analysis in this section is consistent with our main experimental setup in Section 4.1.3. Additional experiments on other Prompts and LLMs are provided in the Appendix A.1.
2.2.1. Empirical Evidence Demonstrating Context Bias
We conduct Feature Ablation Attribution (FAA) (Kokhlikyan et al., 2020; Miglani et al., 2023) analysis to quantify contribution of different token types to the model predictions. FAA is an attribution method used to evaluate the importance of individual input components by measuring how the model’s output changes when specific inputs are masked, and it has been widely adopted in the LLMs for interpreting token-level contributions (Zhou et al., 2024; Zhao et al., 2024; Barkan et al., 2024). A higher attribution value indicates a greater influence of input tokens on the model’s output. Figure 2 shows the ratio of attribution values between auxiliary tokens (e.g., task descriptions or prefix tokens) and user-specific interaction tokens. We report this ratio both before and after fine-tuning to enable direct comparison. From these results, we make the following observation:
Context Bias: Supervised fine-tuning (SFT) can bias LLMs to over-rely on auxiliary tokens while under-utilizing core interaction tokens that encode user personalized preferences.
Before fine-tuning, the attribution value ratio between task description and interaction tokens is approximately 1:1. After fine-tuning, this ratio is markedly amplified across all datasets (e.g., Toy: 4.69:1; Clothing: 6.58:1). A similar phenomenon is observed when comparing prefix tokens with interaction tokens. These results clearly indicate the presence of context bias, whereby the model disproportionately relies on auxiliary tokens rather than more informative interaction tokens.
Figure 4. Proportion of Top‑1 recommendations belonging to Item Group 1 (highest relevance with auxiliary tokens) over the course of SFT.
2.2.2. Negative Effect of Context Bias
Context bias can substantially hinder the effectiveness of LLM-based recommenders, leading not only to decreased recommendation accuracy but also to pronounced unfairness issues. On the one hand, critical user–item interaction signals that capture user preferences may be ignored by the model, severely impairing its ability to deliver personalized recommendations. On the other hand, the recommendation output becomes inherently skewed toward a limited subset of items whose textual tokens exhibit strong correlations with auxiliary tokens.
To empirically verify this phenomenon, we conduct experiments on the typical Amazon datasets. Specifically, we partition items into five groups based on their semantic relevance to auxiliary tokens, measured by the model’s estimated probability:
(2) $r(y)$ $=\frac{1}{|y|}\sum_{t=1}^{|y|}\log P_{\theta}(y_{t}|x^{\text{task}},y_{<t})$
Here we mask the user historical information in the original prompt template with the placeholder ’N/A’, and compute the probability of the tokens for each item. This measure serves as an indicator of the relevance between the input and output, a practice commonly adopted for estimating semantic relevance (Kauf et al., 2024). Based on this metric, items are sorted and evenly divided into five groups, with Group 1 containing the items of highest relevance and Group 5 the least.
Next, we compute the proportion of Top-1 recommended items from each group, with the results shown in Figure 3. The analysis reveals that fine-tuned LLMs display a pronounced inclination for items in Group 1, whose tokens exhibit the highest relevance to auxiliary tokens. Nearly 80% of Top-1 recommendations fall within this group, despite such items comprising only about 20% of the test set. As illustrated in Figure 4, this bias becomes progressively more pronounced as the tuning proceeds. This bias not only undermines recommendation accuracy but also exacerbates exposure unfairness by systematically over-promoting items misaligned with individual user preferences. Such skew can severely degrade user experience and distort the recommendation ecosystem. For example, it would incentivize content providers to adopt clickbait-like titles or other superficial textual strategies to artificially increase token relevance to the auxiliary tokens.
Figure 5. The co-occurrence rate of different types of token pairs in the training set.
2.2.3. Origins of Context Bias
The emergence of context bias can be traced to biases inherent in the training data. As shown in Figure 5, the co-occurrence rate between auxiliary tokens and target item tokens is significantly higher than those between interaction tokens and target item tokens. This is expected, as task prompts appear in every training instance, and prefix tokens are always accompanied for item tokens. Consequently, during fine-tuning, the LLM tends to capture shortcut patterns, simply memorizing frequent correlations with auxiliary tokens, while neglecting the more important user-specific interaction signals. This incurs context bias and motivates the need for improved fine-tuning strategies.
2.2.4. Limitations of Related Strategies
Several recent studies have explored bias and fairness issues in LLM-based RS. However, the characteristics of these biases differ from those of context bias. Specifically, context bias arises during fine‑tuning and reflects the inherent over‑reliance of LLMs on auxiliary tokens. In contrast, earlier work has primarily focused on biases such as popularity bias (Gao et al., 2025b; Lu et al., 2025; Jiang et al., 2024; Gao et al., 2025a; Liao et al., 2024a), arising from imbalanced item frequency in training data; position bias (Chen et al., 2023; Ma et al., 2023; Hou et al., 2024; Bito et al., 2025; Jiang et al., 2025; Dai et al., 2024; Luo et al., 2025; Zhang et al., 2024a), caused by the model’s sensitivity to the ordering of candidate items; amplification bias (Bao et al., 2024), arising from length normalization, which favors items containing tokens with generation probabilities close to 1. These biases stem from distinct sources and mechanisms. Consequently, this newly identified form of context bias warrants explicit and targeted mitigation, yet recent work has fallen short in addressing this issue (cf.  Table 2).
Another related work, CFT (Zhang et al., 2024b), aims to encourage LLMs to better leverage user interaction information. CFT introduces counterfactual learning to forcibly enhance the influence of interaction tokens for each training instance. However, this strategy suffers from multiple aspects of limitations: (1) Objective misalignment. The counterfactual objective is not directly aligned with improving recommendation accuracy or fairness, and its contribution to these aspects remains uncertain. In fact, CFT only treats this objective as an auxiliary loss and over-emphasis on this term has been observed to lead to substantial performance degradation. (2) Difficulty in weight selection. CFT relies on manually specified weights to determine the degree to which a training instance should rely on interaction tokens. In fact, the optimal weights can vary substantially across instances and are difficult to estimate accurately. Manual specification is therefore challenging, prone to deviation from the ideal value, and ultimately detrimental to model effectiveness. Empirically, even when we directly employ the official source code of CFT and conduct a fine‑grained hyperparameter search, CFT yields only limited performance gains on some datasets (cf.  Table 2). (3) High computational overhead. CFT requires to process counterfactual instances, resulting in more than double the training time compared with SFT (cf. Figure 12). In contrast, our GDRT is explicitly aligned with the target objective, enabling the model to consistently perform well across different instance groups. This naturally encourages greater focus on interaction tokens without requiring manual specification of influence strengths.
3. Methodology
The above analyses reveal a significant context bias inherent in fine-tuning LLMs for recommendation. To address this issue, we propose a novel fine‑tuning framework, termed GDRT, which leverages Group Distributionally Robust Optimization (Group DRO) (Sagawa et al., 2019) to mitigate such bias. This section first outlines the general idea of GDRT, and then describes the group partitioning strategy and customized loss function.
General Idea. Rather than directly intervening the learning process to manually enhance the effect of interaction tokens, we pursue an alternative objective: ensuring that the model achieves consistently strong performance across target tokens regardless of their degree of relevance to auxiliary tokens. It naturally shifts the LLM’s attention away from auxiliary tokens towards user‑specific interaction tokens, as simple reliance on shortcut auxiliary tokens results in poor performance on the instances with weak correlations. Besides, this objective is directly aligned with the goals of high recommendation accuracy and fairness.
To implement this idea, we adopt Group DRO, which partitions the training data into multiple groups and uses an adversarial training mechanism to encourage great performance across all groups. Group DRO has been widely applied in various domains and shown effectiveness in mitigating group disparities and shortcut correlations (Sagawa et al., 2019; Zhao et al., 2023; Oren et al., 2019). In applying Group DRO to our problem, we address two key questions: (1) how to construct groups that capture varying degrees of relevance to auxiliary tokens; and (2) how to design a loss function that is computationally efficient.
Token Grouping by Relevance with Auxiliary Tokens. Our objective is to ensure consistent model performance across groups that differ in their degree of relevance to auxiliary tokens. Accordingly, the constructed groupings should capture variations in this relevance. Towards this end, we compute the predictive probability of each target token conditioned solely on the corresponding auxiliary tokens:
(3) $r(y^{*}_{i,t})=\log P_{\theta}\left(y^{*}_{i,t}\mid x^{\text{task}},y^{*}_{i,<t}\right).$
where $y^{*}_{i,t}$ denotes the $t$-th token of the target item in the $i$-th training sample, and $y^{*}_{i,<t}$ denotes its prefix tokens. Based on these relevance scores, all target tokens are partitioned into $G$ disjoint groups $\{\mathcal{G}_{1},\dots,\mathcal{G}_{G}\}$ using the K-means algorithm [^2] We can simply employ the scikit-learn package to implement the K-means algorithm. Considering that $r(y^{*}_{i,t})$ is a numerical value, it can be evenly partitioned according to its magnitude, yielding comparable results., where $G$ is a hyperparameter that determines the granularity of grouping. Each group $\mathcal{G}_{g}$ contains the index $(i,t)$ of target tokens exhibiting similar degrees of relevance to auxiliary tokens, whereas tokens belonging to different groups demonstrate distinct relevance levels.
Unlike Eq. 2, which evaluates item-wise relevance between the entire target item and the auxiliary tokens, Eq. 3 computes token-wise relevance. This distinction is crucial, as different tokens within the same target item may exhibit varying degrees of relevance to the auxiliary tokens. Such finer-grained grouping can more effectively capture variations in this relevance.
Figure 6. Loss of token groups with varying degrees of relevance to auxiliary tokens in the SFT-trained biased model. Target tokens are grouped into strong- and weak-relevance sets using K-means on relevance scores computed with Eq. 3.
The Objective of GDRT. Subsequently, we employ Group DRO to enforce consistent model performance across groups with different strength of relevance. Specifically, the training objective of GDRT is defined as:
(4) $\mathcal{L}_{GDRT}=\max_{Q}\sum_{g=1}^{G}Q(g)\mathcal{L}(g)\quad\text{s.t.}\quad D_{KL}({Q},{U})\leq\eta$
$\mathcal{L}(g)=\frac{1}{|\mathcal{G}_{g}|}\sum_{(i,t)\in\mathcal{G}_{g}}-\log P_{\theta}\left(y^{*}_{i,t}\mid x_{i},y^{*}_{i,<t}\right)$
where $\mathcal{L}(g)$ represents the vanilla generative loss for group $\mathcal{G}_{g}$, and $|\mathcal{G}_{g}|$ represents the number of target tokens in group $\mathcal{G}_{g}$. The term $Q$ denotes the weight distribution over groups, acting as a flexible adversarial perturbation to the original empirical distribution, with $Q(g)$ being the weight assigned to the $g$-th group. This weight distribution is regularized via a Kullback–Leibler divergence term $D_{KL}(Q,U)$ between the weight distribution $Q$ and the uniform distribution $U$, with the parameter $\eta$ controlling the perturbation magnitude.
Comparing $\mathcal{L}_{GDRT}$ (Eq. LABEL:eq:GDRT_loss) with the original SFT objective $\mathcal{L}_{SFT}$ (Eq. 1), the main difference lies in the introduction of the additional group weighting term $Q(g)$. Intuitively, Group DRO imposes an adversarial shift to the group distribution by perturbing the relevance distribution with respect to auxiliary tokens, thereby compelling the model to perform consistently well across groups under different adversarial re‑weightings. This mechanism naturally mitigates the model’s reliance on auxiliary tokens, and encourages the model to perform consistently well across different groups.
Efficient Implementation. While Group DRO involves the complex adversarial optimization, it can be simplified to an equivalent closed‑form objective. We have the following lemma:
Lemma 0.
Equation LABEL:eq:GDRT_loss can be reformulated as the following objective:
(5) $\mathcal{L}_{GDRT}=\sum_{g=1}^{G}Q(g)\mathcal{L}(g),\quad Q(g)=\frac{e^{\mathcal{L}(g)/\tau}}{\sum_{g^{\prime}=1}^{G}e^{\mathcal{L}(g^{\prime})/\tau}}$
The parameter $\tau$ is the dual Lagrange coefficient associated with the constraint $D_{KL}(Q,U)\leq\eta$.
The lemma yields an explicit closed‑form solution for the group weights $Q(g)$. From the perspective of these weights, the effect of Group DRO can be well understood: groups with higher losses, which typically correspond to tokens exhibit weaker relevance with auxiliary tokens (see Figure 6), receive larger weights. This adaptive reweighting encourages the model to allocate greater optimization emphasis to underperforming groups, thereby enhancing its learning on samples with low auxiliary-token relevance. As a result, the model attains more balanced performance across groups and effectively mitigates its over-reliance on auxiliary tokens.
This closed‑form also facilitates an efficient implementation of GDRT. Compared to SFT, the additional steps just involve: partitioning the token groups based on Eq. 3, and dynamically updating the group weights according to Eq. 5. The time complexity of GDRT is the same as the basic SFT, and our empirical experiments also demonstrate their close running time (cf.  Figure 12). Besides, this simple reformulation makes the integration of GDRT into existing LLM-based RS straightforward, requiring only minimal code changes.
Table 1. Statistics of the datasets.
Dataset #Users #Items #Interactions #Density
Toy 19124 11758 165247 0.0735%
Clothing 39230 22948 277534 0.0308%
Book 16559 6344 151928 0.1446%
Table 2. The performance comparison on three real-world datasets. The best result is bolded. Lower MGU and DGU indicate better fairness.
[…]
Table 3. Performance comparison under different LLM-based RS. The best result is bolded.
Method Toy Clothing
NDCG@5 DGU@5 NDCG@5 DGU@5
MSL 0.0198 0.7162 0.0077 0.5815
MSL+GDRT 0.0253 0.2697 0.0096 0.1469
LLaRA 0.0131 0.6485 0.0041 0.5412
LLaRA+GDRT 0.0168 0.4249 0.0053 0.2320
A-LLM 0.0129 0.6122 0.0045 0.5238
A-LLM+GDRT 0.0160 0.4192 0.0053 0.1817
4. Experiments
We aim to answer the following research questions:
  • RQ1: How does GDRT perform compared to existing SOTA methods?
  • RQ2: Can GDRT be integrated into other LLM-based RS?
  • RQ3: Does GDRT mitigate context bias?
  • RQ4: How do the hyperparameters affect the performance of GDRT?
  • RQ5: How does the efficiency of GDRT compare with baseline methods?
4.1. Experimental Settings
4.1.1. Datasets
Three widely used real-world datasets—Amazon Toys and Games, Amazon Clothing, Shoes and Jewelry, and Amazon Books [^3] https://jmcauley.ucsd.edu/data/amazon/index_2014.html—are employed in our experiments. These datasets are commonly adopted in studies on LLM-based recommenders (Cui et al., 2024; Bao et al., 2025; Cao et al., 2024; Lee et al., 2024; Li et al., 2023a; Wang et al., 2025b; Lee et al., 2024). To ensure fair comparison, we follow the data preprocessing procedures employed in recent literature (Bao et al., 2025; Cui et al., 2024; Wang et al., 2025b; Bao et al., 2024). Specifically, we first apply the 5-core setting to the raw datasets. For user interaction sequences exceeding 11 interactions, we segment the sequences using a sliding window of length 11. The segmented sequences are then sorted in ascending order by timestamp and partitioned into training, validation, and test sets in an 8:1:1 ratio. Due to the large size of Amazon Books, we randomly retain 100,000 items prior to 5-core processing. The statistics of the processed datasets are summarized in Table 1.
4.1.2. Baselines
The methods compared fall into several categories:
  • Traditional RS (SASRec (Kang and McAuley, 2018), SASRec++ (Lepage et al., 2025), DROS (Yang et al., 2023)): SASRec employs a self-attention-based model to effectively capture user interests. SASRec++ further improves performance using the softmax loss. DROS integrates DRO to enhance the model’s robustness against distributional shifts.
  • LLM-based RS (SFT (Bao et al., 2025), CFT (Zhang et al., 2024b), MSL (Wang et al., 2025b), LLaRA (Liao et al., 2024b), A-LLM (Kim et al., 2024)): SFT introduces instruction-tuning templates to fine-tune LLMs on RS datasets. CFT incorporates a causal loss to strengthen the behavior sequence modeling capabilities of LLMs. MSL improves the loss function specifically tailoring it to optimize recommendation-oriented objectives. LLaRA incorporates embeddings from traditional models to enhance the ability of LLMs to capture collaborative signals. A-LLM further aligns these embeddings with their corresponding textual information.
  • Debiasing Methods for LLM-based RS (Reweight (Jiang et al., 2024), SPRec (Gao et al., 2025a), D3 (Bao et al., 2024)): Reweight addresses popularity bias by balancing recommendations using pre-calculated item weights. SPRec develops negative sampling techniques based on Direct Preference Optimization (DPO) (Rafailov et al., 2024) to reduce popularity bias. D3 mitigates amplification bias through enhanced decoding strategies.
4.1.3. Implementation Details
For all LLM-based methods, we adopt LLaMA3.2-3B (Dubey et al., 2024) as the backbone, with the number of training epochs set to 5. The prompt design follows (Bao et al., 2025; Wang et al., 2025b). For inference, we employ Constrained Beam Search (CBS), following prior work (Wang et al., 2025b; Bao et al., 2024), to ensure that the recommended items are drawn from the item set. The beam size is fixed at 10. For evaluation, we evaluate the NDCG@5 on the validation set for each epoch’s checkpoint and select the checkpoint with the highest score for testing. The corresponding results on the test set are reported as the final performance. For the hyperparameters in GDRT, the number of groups $G$ is tuned from $\{2,5,10\}$, and $\tau$ is tuned from $\{0.1,0.2,0.3,0.5,1.0\}$. To ensure fair comparisons, we utilize the source code provided by the original authors and tune the hyperparameters of all baseline methods according to the guidelines specified in their respective publications.
Figure 7. Ratio of attribution values between auxiliary tokens and user interaction tokens using SFT and GDRT. Left: task description vs. user interaction tokens. Right: prefix tokens of predicted item (take the first token) vs. user interaction tokens.
4.1.4. Metrics
Four widely used evaluation metrics are employed in this study: NDCG@K and Hit Ratio@K are used to evaluate recommendation accuracy, while MGU@K and DGU@K are adopted to evaluate fairness (Jiang et al., 2024) (K=5, 10). Specifically, for the fairness metrics, items are divided into five groups according to Equation 2. We then calculate the discrepancy between each group’s proportion in the Top-K recommendations and its proportion in the user’s interaction history. MGU@K measures the average of these discrepancies, whereas DGU@K quantifies the gap between the maximum and minimum discrepancies across groups. Accordingly, smaller values of MGU@K and DGU@K indicate that the distribution of recommended items is more aligned with that of the user’s historical interactions, reflecting better fairness.
4.2. Performance Comparison (RQ1 & RQ2)
Table 2 presents a comparative analysis of the proposed GDRT method against the baselines. Overall, GDRT demonstrates substantial performance improvements across all datasets. By mitigating context bias, reducing excessive reliance on auxiliary tokens, and enhancing the model’s ability to capture user-specific behavioral patterns, GDRT delivers marked gains in both accuracy and fairness. Specifically, GDRT achieves an average improvement of 24.29% in NDCG@10 compared with SOTA methods. In terms of fairness, GDRT significantly alleviates the model’s tendency to recommend items strongly correlated with auxiliary information, leading to an average reduction of 37.43% in MGU@5 relative to the SOTA baselines. In contrast, other comparison methods yield only marginal improvements, and their lack of explicit mechanisms to address context bias results in limited fairness enhancement.
Furthermore, we evaluate the performance of GDRT when integrated into other advanced LLM-based RS. As shown in Table 3, GDRT consistently achieves significant performance gains and improves fairness across all methods, indicating that it can be seamlessly incorporated into diverse LLM-based recommenders while exhibiting strong generalization capability.
Figure 8. Distribution of Top‑1 recommended items from GDRT-trained LLMs across five groups defined by auxiliary-token relevance (Group 1: highest relevance, Group 5: lowest relevance), with test-set distribution for comparison.
Figure 9. Proportion of Top‑1 recommendations belonging to Item Group 1 (highest relevance with auxiliary tokens) during training with SFT and GDRT.
4.3. In-depth Analysis (RQ3)
In this section, we present an empirical analysis of the effectiveness of GDRT in mitigating context bias. As shown in Figure 7, GDRT considerably reduces the ratio of attribution values assigned to auxiliary tokens relative to interaction tokens, compared with SFT, indicating that GDRT effectively alleviates the problem of over-reliance on auxiliary tokens. Furthermore, GDRT markedly mitigates the unfairness induced by context bias. As shown in Figure 8, the recommendation distributions generated by GDRT more closely align with the item distribution in the test set. In addition, Figure 9 reveals that GDRT alleviates the issue observed in SFT, where the training process continuously amplifies the recommendation of items highly correlated with auxiliary tokens.
4.4. Hyper-Parameter Sensitivities (RQ4)
In this section, we perform a sensitivity analysis on the hyperparameters of GDRT: the number of groups $G$ in K-means (Figure 10) and the coefficient $\tau$ in DRO (Figure 11). For both parameters, model performance exhibits a trend of initially increasing and then decreasing as the parameter values rise. This phenomenon can be explained as follows. For $G$, an excessively small value may group highly heterogeneous data together, thereby diminishing the differences in auxiliary-token relevance between groups. Conversely, an overly large $G$ results in groups containing too few samples to reliably represent the underlying distribution. As for $\tau$, a smaller value places greater emphasis on groups with higher losses; however, overemphasizing the worst-performing group can cause overfitting and impair overall generalization. In contrast, a larger $\tau$ balances optimization across all groups, but may reduce group-wise consistency in performance. Besides, the model demonstrates strong performance across a wide range of parameter settings, indicating the robustness of GDRT to hyperparameter selection.
Figure 10. Hyperparameter sensitivity analysis on group number $G$ (dashed: SFT baseline).
Figure 11. Hyperparameter sensitivity analysis on parameter $\tau$ in DRO (dashed: SFT baseline).
4.5. Efficiency Comparison (RQ5)
This section compares the efficiency and performance of GDRT with baseline methods. As illustrated in Figure 12, GDRT demonstrates both optimal performance and high computational efficiency. In contrast, the DPO-based method SPRec requires an additional reference model and incorporates extra negative samples during training, leading to a considerable increase in runtime. CFT incurs higher computational costs by introducing additional counterfactual samples during training. Although Reweight and D3 incur negligible training overhead, their performance improvements remain limited.
5. Related Work
5.1. Sequential Recommendation
Sequential recommendation focuses on predicting the next item a user will be interested in based on their historical interactions. Compared to collaborative filtering (Chen et al., 2025a, 2024a; Yang et al., 2024; Zhang et al., 2025a; Yang et al., 2025; Wang et al., 2024a), sequential recommendation incorporates temporal information and places greater emphasis on capturing the evolving patterns of user interests. With the advancement of deep learning, numerous architectures based on deep neural networks have been introduced into the sequential recommendation. For example, GRU4Rec (Hidasi et al., 2015) employs RNNs, while Caser (Tang and Wang, 2018) utilizes CNNs to effectively capture long-term dependencies and modeling user interest patterns from historical behavior. More advanced models such as SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) leverage self-attention mechanisms (Vaswani, 2017), enabling the identification of the most relevant parts within the sequence. Due to the dynamic nature of data distributions as time evolves (Wang et al., 2024c, d), DROS (Yang et al., 2023) introduces DRO (Rahimian and Mehrotra, 2019; Wu et al., 2023a) to further enhance the model’s robustness against distributional shifts caused by temporal changes. The readers may refer to the survey (Fang et al., 2020; Wang et al., 2019) for more details.
Figure 12. Performance comparisons in terms of both recommendation accuracy and efficiency.
5.2. Biases in LLM-based Recommendation
Large Language Models (LLMs), with their powerful capabilities in comprehension, reasoning, and extensive knowledge (Dubey et al., 2024; Achiam et al., 2023; Guo et al., 2025; Team et al., 2024), have been widely applied to recommendation systems (Wu et al., 2024; Wang et al., 2025c; Cui et al., 2025b, a). One prominent paradigm is LLM-based RS (Li et al., 2023c), which directly leverages LLMs as the backbone of the recommender. Subsequent studies have explored fine-tuning LLMs on domain-specific recommendation datasets to enhance their recommendation capabilities (Li et al., 2023c; Bao et al., 2025, 2023; Zhu et al., 2024; Wang et al., 2025b, 2024b; Liao et al., 2024b; Kim et al., 2024; Chen et al., 2024b).
Recent studies have extensively explored bias and fairness issues in LLM-based RS, such as popularity bias (Gao et al., 2025b; Lu et al., 2025; Jiang et al., 2024; Gao et al., 2025a; Liao et al., 2024a; Lin et al., 2025), position bias (Chen et al., 2023; Ma et al., 2023; Hou et al., 2024; Bito et al., 2025; Jiang et al., 2025; Dai et al., 2024; Luo et al., 2025; Zhang et al., 2024a), amplification bias (Bao et al., 2024), and bias stemming from LLMs’ preferences for specific item attributes (Li et al., 2023b; Deldjoo, 2024; Jiang et al., 2024; Zhang et al., 2025b). Nevertheless, existing research has largely overlooked context bias, which arises during fine-tuning and reflects the inherent over-reliance of LLMs on auxiliary tokens. Since existing debiasing methods fail to account for this factor, their effectiveness remains limited. CFT (Zhang et al., 2024b) seeks to enhance the modeling of users’ historical interactions, but suffers from objective misalignment, weight selection challenges, and high computational cost, restricting its applicability. These limitations are further discussed in Section 2.2.4.
5.3. Group Distributionally Robust Optimization
Group Distributionally Robust Optimization (Group DRO) (Sagawa et al., 2019) is an optimization framework that operates over predefined sample groups, aiming to achieve consistent and reliable performance across them by emphasizing the optimizing of the worst-performing group during training. It has been widely applied in various domains (Sagawa et al., 2019; Zhao et al., 2023; Oren et al., 2019) and has demonstrated strong effectiveness in mitigating group disparities (Hashimoto et al., 2018; Zhao et al., 2023) as well as reducing models’ reliance on shortcut correlations (Han et al., 2024; Creager et al., 2021). Several studies have also integrated Group DRO into RS to improve model performance for sparse users and items (Wen et al., 2022; Zhao et al., 2023).
6. Conclusion
In this work, we identify a key limitation of supervised fine-tuning (SFT) in LLM-based recommenders: it often induces Context Bias, whereby the model over-relies on auxiliary tokens (e.g., task descriptions and prefix-generated tokens) while underutilizing core user interaction information. This bias undermines recommendation accuracy and raises unfairness concerns. To address this issue, we introduce Group Distributionally Robust Optimization-based Tuning (GDRT), which aims to reduce the model’s over-reliance on auxiliary tokens by applying Group DRO across token groups with varying degrees of relevance to auxiliary tokens. Extensive experiments on multiple public datasets demonstrate that GDRT effectively mitigates context bias, thereby significantly improving recommendation accuracy and enhancing fairness.
This work investigates a novel form of bias introduced by the integration of LLMs into recommenders, which is not present in traditional recommendation models. A promising avenue for future work is to examine whether LLM-based RS exhibit additional, as-yet unidentified biases.
Appendix A Appendices
A.1. Verification of Context Bias Across Diverse Prompt Templates and LLMs
In this section, we extend the FAA experiments described in Section 2.2.1 to a broader range of prompt templates (see Figure 13) and LLM backbones (see Figure 14), in order to assess whether context bias is a widespread phenomenon. For prompt templates, we adopt those proposed in (Liao et al., 2024b) and (Bao et al., 2024), referred to as Prompt1 and Prompt2. For LLM backbones, we evaluate LLaMA3‑8B (Dubey et al., 2024) and Qwen2.5‑1.5B (Team et al., 2024). Across all prompt and backbone configurations, the results consistently reveal pronounced context bias, wherein SFT leads the model to over-rely on auxiliary tokens.
Figure 13. Ratio of attribution values between auxiliary tokens and user-interaction tokens before and after SFT across different prompt templates.
Figure 14. Ratio of attribution values between auxiliary tokens and user-interaction tokens before and after SFT across different LLMs.
A.2. The proof of Lemma 3.1
The original formulation of GDRT
(6) $\mathcal{L}_{GDRT}=\max_{Q}\sum_{g=1}^{G}Q(g)\mathcal{L}(g),\quad\text{s.t. }D_{KL}(Q,U)\leq\eta,$
can be rewritten in the expectation form as
(7) $\max_{Q}\mathbb{E}_{g\sim Q}[\mathcal{L}(g)]$
$s.t.$ $\mathbb{E}_{g\sim Q}[\log\frac{Q(g)}{U(g)}]\leq\eta$
In the following, we focus on how to eliminate the inner maximization optimization problem and the KL constraint term. Assume $W(g)=Q(g)/U(g)$ and define a convex function $\phi(x)=x\log x-x+1$. Then the divergence $D_{KL}({Q},{U})$ can be written as $\mathbb{E}_{U}[\phi(W)]$. The inner layer maximization optimization problem can be reformulated as follow:
(8) $\max_{W}\mathbb{E}_{U}\left[\mathcal{L}W\right]$
s.t. $\mathbb{E}_{U}[\phi(W)]\leq\eta,\mathbb{E}_{U}[W]=1$
As a convex optimization problem, we use the Lagrangian function to solve it:
(9) $\min_{\tau\geq 0,\beta}\max_{W}\mathbb{E}_{U}\left[\mathcal{L}W\right]-\tau(\mathbb{E}_{U}[\phi(W)]-\eta)+\beta(\mathbb{E}_{U}[W]-1)$
$=\min_{\tau\geq 0,\beta}\left\{\tau\eta-\beta+\tau\max_{W}{\mathbb{E}_{U}}\left[\frac{\mathcal{L}+\beta}{\tau}W-\phi(W)\right]\right\}$
$=\min_{\tau\geq 0,\beta}\left\{\tau\eta-\beta+\tau\mathbb{E}_{U}\left[\max_{W}\left(\frac{\mathcal{L}+\beta}{\tau}W-\phi(W)\right)\right]\right\}$
Notice that $\max_{W}\left(\frac{\mathcal{L}+\beta}{\tau}W-\phi(W)\right)=\phi^{*}(\frac{\mathcal{L}+\beta}{\tau})$ is the convex conjugate function of $\phi(x)$ and we have $\phi^{*}(x)=e^{x}-1$. $W(g)=e^{\frac{\mathcal{L}(g)+\beta}{\tau}}$ when the maximum value is obtained.
(10) $\min_{\tau\geq 0,\beta}\left\{\tau\eta-\beta+\tau\mathbb{E}_{U}\left[\max_{W}\left(\frac{\mathcal{L}+\beta}{\tau}W-\phi(W)\right)\right]\right\}$
$=\min_{\tau\geq 0,\beta}\left\{\tau\eta-\beta+\tau\mathbb{E}_{U}\left[e^{\frac{\mathcal{L}+\beta}{\tau}}-1\right]\right\}$
$=\min_{\tau\geq 0}\left\{\tau\eta+\tau\log{\mathbb{E}_{U}\left[e^{\frac{\mathcal{L}}{\tau}}\right]}\right\}$
where $\beta=-\tau\log{\mathbb{E}_{g\sim U}\left[e^{\frac{\mathcal{L}(g)}{\tau}}\right]}$ and $W(g)=\frac{e^{\frac{\mathcal{L}(g)}{\tau}}}{\mathbb{E}_{g^{\prime}\sim U}\left[e^{\frac{\mathcal{L}(g^{\prime})}{\tau}}\right]}$ when the minimum value is obtained. We consider the Lagrange multiplier $\tau$ as a hyperparameter related to the robustness radius $\eta$. Then we can get the unconstrained optimization problem as follows,
(11) $\mathcal{L}_{GDRT}=\tau\eta+\tau\log\mathbb{E}_{g\sim U}\mathrm{exp}\left(\frac{\mathcal{L}(g)}{\tau}\right)$
where the worst-case distribution
(12) $Q^{*}(g)=U(g)\frac{\mathrm{exp}\left(\mathcal{L}(g)/\tau\right)}{\mathbb{E}_{i\sim U}\left[\mathrm{exp}(\mathcal{L}(i)/\tau)\right]}$
Since $U$ is a uniform distribution
(13) $Q^{*}(g)=\frac{\mathrm{exp}\left(\mathcal{L}(g)/\tau\right)}{\sum_{g^{\prime}}\left[\mathrm{exp}(\mathcal{L}(g^{\prime})/\tau)\right]}$
Thus lemma 3.1 is proven.